40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9877     Entry time: Wed Apr 30 00:40:55 2014
Author: manasa 
Type: Configuration 
Category: LSC 
Subject: Y arm IR lock troubleshooting 

[Koji, Manasa]

The Y arm locks stably for IR PDH now.  

The reason for ETMY getting kicked during lock acquisition was finally found to be related to the limiter value set in the Y arm servo. We reduced the limiter value unintentionally and found that the lock acquisition stayed smooth. The limiter value was stepped in 1000s from 7000 and eventually found that the ETMY suspension was kicked when we try to acquire lock with the limiter value was set at 11000. 

 

The limiter for X arm at 11000 is not causing any trouble at the moment.

In the process, we did a bunch of things through the evening to troubleshoot IR locking of the Y arm.

Earlier today running the IFO configure script did not restore the arms to lock and both the ETMs needed to be aligned to lock the arms. The arms stayed locked for 15 minutes and the Y arm lost lock eventually leaving the ETMY in a misaligned state. 

The state of the Y arm was similar to what Jenne has explained in ELOG where the ETMY was kicked during lock acquisition and would move to a misaligned state.

To trouble shoot, there were several things that were done. A few of them might not have any direct correlation to the locking issue but could just be a coincidence.  

1.  The trigger time for the filters in the arm filter modules were set such that they switch on after the SUS violin filters. Arm FM trigger time = 3 s (previously set at 0.1s) and SUS violin trigger time = 1s. This reduced the number of lock loss events.

2. There was some drop in transmission when the bounce filter of Y arm (FM6) turned ON. This was fixed by changing its ramp time (initially set at 1s). The filter has been modified to turn on immediately upon arm lock acquisition before the other triggered filters in the filter module turn on.

3. The QPD and SUS signal cables running to the rack were checked to be intact. Koji found some of them to be loose. But this had no evident correlation with the arm locking problem.

4.The oplev and PD alignment was checked at the Y end. The high gain trans PD for Y arm was checked for good alignment by looking at TRY. It was found that the EXIT light at the Y end is injecting some noise to the transmission PD. 

5. The ETMY was given offsets in PIT, YAW and POS and the OSEM sensor values were checked to see if the suspension is behaving well. It was behaving well.

ELOG V3.1.3-